e Period ( Radius (m) Velocity (m/s) Acceleration (m/) 12T 96π 144π 729π A. 9 m
Answer #1

Circle A :

Period T = 12 \pi

we have equation T = 2\pi / \omega

2\pi / \omega = 12 \pi

\omega = 0.167 rad/sec

using \omega = 2\pir

r = \omega / 2\pi

= 0.167 / 2\pi

r = 0.0265 m

using V = r \omega

V = 0.0265 x 0.167

V = 0.004425 m/sec

using a = V2 / r

= 0.0044252 / 0.0265

a = 0.0007388 m/sec2

Circle B :

Period T = 96 \pi

we have equation T = 2\pi / \omega

2\pi / \omega = 96 \pi

\omega = 0.021 rad/sec

using \omega = 2\pir

r = \omega / 2\pi

= 0.021 / 2\pi

r = 0.0033 m

using V = r \omega

V = 0.0033x 0.021

V = 0.000693 m/sec

using a = V2 / r

= 0.0006932 / 0.0033

a = 0.0001455 m/sec2

Circle C :

Period T = 144 \pi

we have equation T = 2\pi / \omega

2\pi / \omega = 144 \pi

\omega = 0.0139 rad/sec

using \omega = 2\pir

r = \omega / 2\pi

= 0.0139 / 2\pi

r = 0.0022 m

using V = r \omega

V = 0.0022 x 0.0139

V = 0.0000306 m/sec

using a = V2 / r

= 0.00003062 / 0.0022

a = 0.000000425 m/sec2

Circle D :

Period T = 729 \pi

we have equation T = 2\pi / \omega

2\pi / \omega = 729 \pi

\omega = 0.0027 rad/sec

using \omega = 2\pir

r = \omega / 2\pi

= 0.0027 / 2\pi

r = 0.00043 m

using V = r \omega

V = 0.00043 x 0.0027

V = 0.00000116 m/sec

using a = V2 / r

= 0.000001162 / 0.00043

a = 0.00000000313 m/sec2